OpenStudy (anonymous):

. Gaseous effusion can be used to separate isotopes. Suppose you had a balloon filled with equal molar amounts of hydrogen, H2, and deuterium, D2, gas.  After 10 hours you find that 20.0 % of the hydrogen, H2, had effused out of the balloon along with some unknown amount of the deuterium, D2.    What would be the molar percentage of D2 gas in the balloon after this 10 hour period? (Atomic Masses:         H   1.01 grams/mol;     D  2.01 grams/mol)     

OpenStudy (aaronq):

use grahams law and find how much of the Deuterium gas effused

OpenStudy (anonymous):

but how do i set it up>

OpenStudy (aaronq):

\[V _{1/}V _{2}= \sqrt{M _{2}/M _{1}}\] v1=v2 after 10 hrs you lost 20% so v1= 80 % = 0.8

OpenStudy (aaronq):

oh wait no

OpenStudy (aaronq):

i haven't used the formula i forgot that v meant velocity.. i was thinking volume

OpenStudy (aaronq):

so 20% in 10 hour .. thats 20%/10 = 2%/hr so v1= 0.02

OpenStudy (anonymous):

so .02/v2 = sqrt of 1/2 which is .02/v2=1/4 so v2=.08 %?

OpenStudy (aaronq):

\[\frac{ 0.02 }{ V _{2} }=\sqrt{\frac{ 4 }{ 2 }}\]

OpenStudy (aaronq):

V2 =0.0141

OpenStudy (anonymous):

where did you get the 4 from?

OpenStudy (aaronq):

Deuterium gas .. is diatomic, just like hydrogen gas. 1 deuterium atom = 2g/mol Deuterium gas 4 g/mol

OpenStudy (anonymous):

how are we supposed to know that?

OpenStudy (aaronq):

know what? that theres 2 molecules in the gas?

OpenStudy (anonymous):

yes, i still dont get why its 4/2

OpenStudy (aaronq):

\[\frac{ V _{1} }{ V _{2} }=\sqrt{\frac{ M _{2} }{ M _{1} }} \] M = molar mass of the gas V = velocity at which gas diffuses notice that the order is inverted underneath the squared root sign M2 = deuterium gas if you set V2 to be the velocity of effusion for deuterium M1= hydrogen gas if you set V1 to be hydrogen's

OpenStudy (aaronq):

You could also write it like this: \[\frac{ V _{2} }{ 0.02 }=\sqrt{\frac{ 2 }{ 4 }}\]

OpenStudy (aaronq):

as the order is arbitrary

OpenStudy (anonymous):

but it says h=1 and d=2 , why do we double it?

OpenStudy (anonymous):

OOOOO I GET IT LOL

OpenStudy (anonymous):

thank you :)

OpenStudy (aaronq):

haha okay good, no prob dude!